OpenStudy (anonymous):

I used the substitution method to solve 10x+y=–1 5x+2y=16 and I got (-6/5, 11)..

OpenStudy (helder_edwin):

so?

OpenStudy (anonymous):

So.. Is it right?

OpenStudy (anonymous):

yes it is right

OpenStudy (helder_edwin):

in (-6/5,11) you have x=-6/5 and y=11. replace these in the equations see what happens. for instance 10(-6/5)+11=-12+11=-1 what do you think?